# Hot Rocks Survey III: A deep eclipse for LHS 1140c and a new Gaussian process method to account for correlated noise in individual pixels

### v1

@@ 2.2 Introduction to our new pixel-fitting method
Gaussian processes are a method often used to fit correlated noise (also known as systematics) in exoplanet light curves (e.g. Gibson et al. 2012, 2013; Evans et al. 2015; Foreman-Mackey et al. 2017; Aigrain & Foreman-Mackey 2023). To apply them to an eclipse light curve, we must first model the eclipse signal using a deterministic transit model (e.g. using equations from Mandel & Agol 2002). This deterministic model is referred to as the mean function , and the underlying assumption of the method is that we can model the noise around this signal as a random draw from a covariance matrix built using a kernel function. That is to say, we model the noise as having a given amplitude or height scale , and the covariance/correlation in the noise between any two points is described by a kernel function which typically decreases as a function of the separation of the two points using some metric such as time. For example, the exponential kernel combined with white noise is a common choice for a kernel function:
where , could be the timestamps of two datapoints, is the height scale and the length scale of correlated noise, is the amplitude of (uncorrelated) white noise and is the Kronecker delta which equals unity when and zero otherwise.
If we want to fit multiple time-series simultaneously - such as joint-fitting individual pixel light curves - then we can use a 2D GP model, where the dataset lies along a 2D grid of pixels-by-time. In this case, our kernel function can describe correlation as a function of separation in time and as a function of distance away on the detector. As the computational cost of GPs can scale as the cube of the total number of datapoints i.e. for pixels and time points, we need to exploit a particular GP optimisation to make this method computationally tractable for our datasets. We chose the optimisation originally identified in Rakitsch et al. 2013, and introduced to exoplanet time-series by Fortune et al. (2024), as it has sufficient flexibility in the choice of kernel functions we can use and scales efficiently as , making the analysis computationally feasible. The method calculates the exact log likelihood value without approximations and works by breaking up the expensive decomposition of the covariance matrix into separate covariance matrices for each dimension, see Fortune et al. (2024) for more details. It is implemented in an open-source Python package called luas which is available on GitHub111https://github.com/markfortune/luas along with documentation and tutorials (Fortune et al. 2024).
@@ 2.3 Pixel-fitting mean function
First we will need to model the PSF and account for how it will move in each integration due to telescope pointing instability. The PSF was not found to be well-characterised by simple parameterisations such as a 2D Gaussian profile or using the models generated by the package webbpsf (Perrin et al. 2014) and so instead we directly fit for the average flux in time for each pixel included in the fit. We also fit for a linear trend in flux for each pixel (with the slope denoted as ) as different pixels can show quite distinct slopes in time throughout the observations. To account for movements in telescope pointing, we fit for shifts in the PSF position along the and directions for each integration. Finally, the amplitude of the PSF as a function of time is fit with an eclipse model with a single shared set of eclipse parameters featuring the eclipse depth , period , system scale , radius ratio , impact parameter , central transit time , eccentricity and argument of periastron . We also included a shared exponential ramp model between all pixels with height scale and length scale in order to account for the sharp changes in flux typically seen in the first minutes of the observations, although this ramp does appear quite different between pixels in the first minutes of the observations (as will be discussed in Section 4.5) so we avoid fitting the first 45 minutes of data with this method. Alternatively it may be possible to fit this with a separate ramp for each pixel but we did not explore this in this work.
@@
where ) is the flux of pixel as calculated by a 2D interpolating function which takes a grid of average flux values located at the central position of each pixel and linearly interpolates to a new grid offset by i.e. the and shift in PSF position for the integration at time . may be thought of as the baseline flux for each pixel light curve, while and fit for the time-series of telescope pointing movements. , and are all fit for as free parameters. Note we choose to fit for and instead of and as it puts a simple uniform prior on (Eastman et al. 2013).
@@ 2.4 Pixel-fitting kernel function
There are multiple noise processes which are correlated between detector pixels which are not normally required to be accounted for when fitting an aperture extracted light curve. However, if we are to fit an eclipse depth directly using the pixel light curves then we must account for the correlation in the noise between different pixels in order to avoid underestimating or overestimating our eclipse depth uncertainty. We will present each noise process to be accounted for in this section.
@@
where we define two pixel kernel functions and and two time kernel functions and . The terms in and will describe systematics which are correlated in time, while terms in and will describe white noise or correlations between pixels which are instantaneous in time. We will introduce each noise process in turn over the rest of this section but the overall form of our kernel function will be as follows:
where we have three time-correlated systematics processes which due to the restrictions of this GP optimisation must all share the same time kernel function with a shared length scale (chosen to be the exponential kernel), but have different correlations as a function of separation of pixels on the detector. These correlations between pixels are denoted by the pixel kernel functions to be described in this section: , and . We also have multiple noise processes which we model as independent in time including white noise which is adjusted to account for interpixel capacitance (IPC) and correlations associated with the background subtraction. These are described by the pixel kernel functions and . Note that our kernel function is in the form of Equation 3 and that none of the pixel kernel functions can have any time-dependence.
Depending on the noise process it can either be simpler to describe the kernel function in terms of the covariance matrices that it builds or by writing out the kernel function explicitly, so we sometimes switch between these notations. Our kernel function restriction may equivalently be expressed with Equation 5, where the full covariance matrix of the dataset must be expressible as the sum of two Kronecker products
where we have defined two pixel covariance matrices and (generated using the kernel functions and ) and two time covariance matrices and (generated using and ). The Kronecker product is an operation where each element in the matrix left of the sign is multiplied by the full matrix on the right. For two matrices of shape and this will result in a matrix of shape . As our choice of GP optimisation only requires calculating the separate pixel and time covariance matrices, we avoid memory issues from building the full covariance matrix. In addition, instead of needing to decompose the full covariance matrix (e.g. using Cholesky factorisation or eigendecomposition), we only need to separately decompose the pixel and time covariance matrices (specifically using eigendecomposition), significantly reducing computational costs.
@@ 2.4.1 Flux-conserved systematics (FCS)
Examining the pixel light curves reveals significant systematics which do not appear obvious in the aperture extracted light curves (see Figure 1). They are particularly strong in the highest flux pixels at the centre of the PSF. This could be interpreted as the PSF continuously changing in time on timescales of minutes, perhaps due to thermal distortion of mirrors on JWST (Rigby et al. 2023) or alternatively may be due to charge migration between pixels. In principle, we could examine whether these systematics arise from a global process (e.g. from distortion of the mirrors) or from a local process (e.g. from charge migration) if we had another bright star on the detector to compare the pixel light curves to (similar to the ”causal pixel model”; Wang et al. 2016). Unfortunately, we could not find an example of this in a MIRI/Imaging time-series dataset. The systematics not being apparent in the aperture extracted light curve suggests that they may be flux-conserving i.e. a bump up in one pixel may be paired with a bump down in a neighbouring pixel. The autocorrelation of neighbouring pixel light curves from our observations hints that this may be true (see Appendix H). Regardless of the underlying cause, we modelled this using anti-correlated Gaussian processes between neighbouring pixels and we fit for the amplitude of this effect. This allows the model to fit these bumps in individual pixel light curves while keeping the overall aperture sum flux-conserved. We can then fit for additional non-flux-conserving systematics which may arise from other causes which we might be able to weight away from.
@@
We used the exponential kernel to describe this correlation as a function of time, and our pixel kernel function to describe how these systematics are (anti-)correlated between pixels. To model the flux as conserved we assumed that the amount of flux ”moving between” two neighbouring pixels and is proportional to the geometric mean of the baseline flux falling on each pixel . We may then model the correlation in flux between both pixels with a covariance matrix as:
@@
This covariance matrix can describe correlated noise which is perfectly anti-correlated between two pixels, meaning this correlated noise will sum to zero if the pixels are summed within an aperture extraction222Note that while this matrix is only positive semi-definite and not invertible, once white noise is included in the overall kernel then the full covariance matrix will be invertible..
@@
Each pixel will actually be surrounded by eight neighbours, four pixels which share an edge and four pixels which share a corner. For simplicity, we might assume that flux movement occurs with equal amplitude in all directions including between adjacent and diagonal pixels. We can describe this by summing the covariance terms corresponding to flux movement between each neighbouring pair of pixels together. Given these assumptions, we can describe the pixel kernel function of flux-conserved systematics for all pixels using:
@@
Note that we will not be fitting all pixels on the detector which means we will be fitting pixels which have neighbours not included in our fit. As flux may still ”move between” these pixels, we will still account for the additional variance experienced by these border pixels by including these pixel fluxes in the sum of neighbouring fluxes . However, there will be no corresponding covariance terms included in the covariance matrix. This has the consequence that the flux which moves between pixels being fit and pixels not being fit will not be conserved, although this only affects pixels at the edges of our boundary which will typically experience low flux. This effect can be reduced by including more pixels in the fit, although this comes at increased computational cost.
@@ 2.4.2 Independent pixel systematics
Each pixel may also have systematics which are independent to neighbouring pixels, due to processes such as cosmic ray hits, unmasked bad pixels, etc. For example, the pixel light curves from the second LHS 1140c eclipse appear to show a jump in flux in only a couple of pixels (as highlighted in Figure 1). For these systematics we will assume they can be modelled with the same kernel function in time as for flux-conserving systematics and have a pixel kernel function given by:
where we refer to these as independent pixel systematics (IPS) with height scale . Since the amplitude of systematics may vary significantly between pixels (e.g. due to persistence effects from a cosmic ray hitting specific pixels), we may choose to fit for the amplitude separately for each pixel. Checking which pixels recover large values of could help identify pixels exhibiting increased systematics.
@@ 2.4.3 Common systematics between all pixels
We may also encounter systematics which have the same shape in all pixels and are proportional to the flux on each pixel. For example, if the target star flares, that could produce the same systematic shape in all pixels. These systematics may therefore represent stellar activity or alternatively could result from variations in the throughput of the detector which are correlated in time. Note that these systematics should be similar to fitting the aperture extracted light curve with a GP as they are not conserved and each pixel is treated similarly. Our pixel kernel function for these systematics is described by:
We refer to these as common systematics (CS) with height scale . The limitations of our GP optimisation will require these systematics to also share the same time kernel function as the flux-conserved systematics and independent pixel systematics. A comparison between the three previous systematics described is shown in Figure 2.
@@ 2.4.4 Interpixel capacitance (IPC)
This effect can be modelled by taking the image we expect to receive (including Poisson noise) and using a 3x3 convolution kernel (not to be confused with a GP kernel function) which blurs the signal received in each pixel into its neighbouring eight pixels. The kernel will always sum to unity which results in no change to the overall flux received. We can parameterise this convolution kernel using Equation 11 with the coupling between neighbouring pixels to the left or right, for pixels above and below and for diagonal pixels. More complex coupling is also possible but this parameterisation is common in the literature (Moore et al. 2006; Morrison et al. 2023).
Morrison et al. (2023) suggest that for the MIRI detector while is significantly smaller yet still non-zero. The exact extent of IPC is difficult to quantify and is claimed to vary between observations so we will rely on fitting for and for each observation.
We will need to account for both the amplitude of white noise for each pixel and the effect of IPC in our combined pixel kernel function . However, as IPC can be modelled using linear transformations, it is more natural to describe this by the covariance matrix generated by this kernel function . Suppose each pixel observes Poisson noise with variance - which in the absence of IPC is independent between pixels. This can be described by a diagonal covariance matrix which contains the variance of each pixel along its diagonal (the pixels must be sorted into some arbitrary order). We can simulate IPC by taking a random draw from and then convolving with our IPC kernel from Equation 11. Note we can construct a matrix which performs this convolution for each pixel via matrix multiplication , where is our random draw of white noise with the effect of IPC simulated. The covariance matrix of is then given by:
@@
2.4.5 Background flux and row/column systematics
@@
where is the amplitude of overall background scatter, is the amplitude of correlated scatter in each row and similarly for each column. By fitting for the amplitude of , and , we can constrain how much remaining background scatter is left by the background subtraction. We could potentially use this to compare different background subtraction procedures. Note that this will not be able to detect whether the background has been oversubtracted or undersubtracted but simply whether there is a lot of correlated scatter between rows, columns or all pixels.
@@ 3 Testing on simulated data
To better examine how the pixel-fitting method compares to aperture photometry, we generated synthetic observations for which the true eclipse depth was known. This then allows us to examine how precisely and robustly each method recovers the eclipse depth in different circumstances. We examined how well each method performed when there were only flux-conserved systematics present, when systematics were added in only a single pixel light curve and when there were systematics present across all pixels with the same shape (e.g. from a flare).

### v2

@@ 2.2 Introduction to our new pixel-fitting method
Gaussian processes are a method often used to fit correlated noise (also known as systematics) in exoplanet light curves (e.g. Gibson et al. 2012, 2013; Evans et al. 2015; Foreman-Mackey et al. 2017; Aigrain & Foreman-Mackey 2023). To apply them to an eclipse light curve, we must first model the eclipse signal using a deterministic transit model (e.g. using equations from Mandel & Agol 2002). This deterministic model is referred to as the ‘mean function’, , and the underlying assumption of the method is that we can model the noise around this signal as a random draw from a covariance matrix built using a kernel function. That is to say, we model the noise as having a given amplitude or height scale, , and the covariance in the noise between any two points is described by a kernel function which typically decreases as a function of the separation of the two points using a specific metric, such as time. For example, the exponential kernel combined with white noise is a common choice for a kernel function:
where , could be the timestamps of two datapoints, is the height scale and the length scale of correlated noise, is the amplitude of (uncorrelated) white noise, and is the Kronecker delta which equals unity when and zero otherwise.
If we want to fit multiple time series simultaneously (e.g. joint-fitting individual pixel light curves), then we can use a 2D GP model, where the dataset lies along a 2D grid of pixels-by-time. In this case, our kernel function can describe correlation as a function of separation in time and as a function of distance away on the detector. As the computational cost of GPs can scale as the cube of the total number of datapoints, namely, for pixels and time points, we need to exploit a particular GP optimisation to make this method computationally tractable for our datasets. We chose the optimisation originally identified in Rakitsch et al. 2013, and introduced to exoplanet time series by Fortune et al. (2024), as it has sufficient flexibility in the choice of kernel functions we can use and it scales efficiently as , making the analysis computationally feasible. The method calculates the exact log likelihood value without approximations and works by breaking up the expensive decomposition of the covariance matrix into separate covariance matrices for each dimension, see Fortune et al. (2024) for more details. It is implemented in an open-source Python package called luas which is available on GitHub111https://github.com/markfortune/luas along with documentation and tutorials (Fortune et al. 2024).
@@ 2.3 Pixel-fitting mean function
First, we need to model the PSF and account for how it will move in each integration due to telescope pointing instability. The PSF was not found to be well-characterised by simple parameterisations such as a 2D Gaussian profile or using the models generated by the package webbpsf (Perrin et al. 2014) and so instead we directly fit for the average flux in time for each pixel included in the fit. We also fit for a linear trend in flux for each pixel (with the slope denoted as ) as different pixels can show quite distinct slopes in time throughout the observations. To account for movements in telescope pointing, we fit for shifts in the PSF position along the and directions for each integration. Finally, the amplitude of the PSF as a function of time is fit with an eclipse model with a single shared set of eclipse parameters featuring the eclipse depth, ; period, ; system scale, ; radius ratio, ; impact parameter, ; central transit time, ; eccentricity, and argument of periastron, . We also included a shared exponential ramp model between all pixels with height scale, , and length scale, , to account for the sharp changes in flux typically seen in the first minutes of the observations; however, this ramp does appear quite different between pixels in the first minutes of the observations (as discussed in Section 4.5). Thus, we avoided fitting the first 45 minutes of data with this method. Alternatively it may be possible to fit this with a separate ramp for each pixel but we did not explore this in this work.
@@
where ) is the flux of pixel as calculated by a 2D interpolating function which takes a grid of average flux values located at the central position of each pixel and linearly interpolates to a new grid offset by namely, the and shift in PSF position for the integration at a time, ; may be thought of as the baseline flux for each pixel light curve, while and fit for the time series of telescope pointing movements. Then, , and were all fit for as free parameters. We note we chose to fit for and instead of and as it puts a simple uniform prior on (Eastman et al. 2013).
@@ 2.4 Pixel-fitting kernel function
There are multiple noise processes which are correlated between detector pixels that are not normally required to be accounted for when fitting an aperture extracted light curve. However, if we are trying to fit an eclipse depth directly using the pixel light curves, then we must account for the correlation in the noise between different pixels in order to avoid underestimating or overestimating our eclipse depth uncertainty. We present each noise process to be accounted for in this section.
@@
Here, we define two pixel kernel functions, and , and two time kernel functions, and . The terms in and describe the systematics which are correlated over time, while the terms in and describe white noise or correlations between pixels which are instantaneous in time. We introduce each noise process in turn over the rest of this section, but the overall form of our kernel function is
where we have three time-correlated systematics processes. Due to the restrictions of this GP optimisation, they must all share the same time kernel function with a shared length scale (chosen to be the exponential kernel), but with different correlations as a function of separation of pixels on the detector. These correlations between pixels are denoted by the pixel kernel functions to be described in this section: , , and . We also have multiple noise processes that we modeled as independent in time, including correlations associated with the background subtraction, as well as white noise which must be adjusted to account for interpixel capacitance (IPC). These are described by the pixel kernel functions and . Note that our kernel function is in the form of Equation 3 and that none of the pixel kernel functions can have any dependence on time.
Depending on the noise process, it can either be simpler to describe the kernel function in terms of the covariance matrices that it builds or by writing out the kernel function explicitly; thus, we sometimes switch between these notations. Our kernel function restriction may equivalently be expressed with Equation 5, where the full covariance matrix of the dataset must be expressible as the sum of two Kronecker products
Here, we have defined two pixel covariance matrices and (generated using the kernel functions and ) and two time covariance matrices and (generated using and ). The Kronecker product is an operation where each element in the matrix left of the sign is multiplied by the full matrix on the right. For two matrices of shape and this will result in a matrix of shape . Since our choice of GP optimisation only requires calculating the separate pixel and time covariance matrices, we avoid memory issues from building the full covariance matrix. In addition, instead of needing to decompose the full covariance matrix (e.g. using Cholesky factorisation or eigendecomposition), we only need to separately decompose the pixel and time covariance matrices (specifically using eigendecomposition), significantly reducing the computational costs.
@@ 2.4.1 Flux-conserved systematics (FCS)
Examining the pixel light curves reveals significant systematics that do not appear obvious in the aperture extracted light curves (see Figure 1). They are particularly strong in the highest flux pixels at the centre of the PSF. This could be interpreted as the PSF continuously changing in time on timescales of minutes, perhaps due to thermal distortion of mirrors on JWST (Rigby et al. 2023) or alternatively may be due to charge migration between pixels. In principle, we could examine whether these systematics arise from a global process (e.g. from distortion of the mirrors) or from a local process (e.g. from charge migration) if we have another bright star on the detector to compare the pixel light curves to (similar to the ‘causal pixel model’; Wang et al. 2016). Unfortunately, we could not find an example of this in a MIRI/Imaging time-series dataset. The systematics not being apparent in the aperture extracted light curve suggests that they may be flux-conserving; namely, a bump up in one pixel may be paired with a bump down in a neighbouring pixel. The autocorrelation of neighbouring pixel light curves from our observations hints that this may be true (see Appendix H). Regardless of the underlying cause, we modelled this using anti-correlated GPs between neighbouring pixels and we fit for the amplitude of this effect. This allows the model to fit these bumps in individual pixel light curves while keeping the overall aperture sum flux-conserved. We can then fit for additional non-flux-conserving systematics, which may arise from other causes that we might be able to weight away from.
@@
We used the exponential kernel to describe this correlation as a function of time and our pixel kernel function to describe how these systematics are (anti-)correlated between pixels. To model the flux as conserved we assumed that the amount of flux ‘moving between’ two neighbouring pixels, and , is proportional to the geometric mean of the baseline flux falling on each pixel, . We can then model the correlation in flux between both pixels with a covariance matrix as:
@@
This covariance matrix can describe correlated noise which is perfectly anti-correlated between two pixels, meaning that the sum of this correlated noise will be zero if the pixels are summed within an aperture extraction222Note: while this matrix is only positive semi-definite and not invertible, once white noise is included in the overall kernel then the full covariance matrix will be invertible..
@@
Each pixel actually ends up surrounded by eight neighbours, four pixels which share an edge and four pixels which share a corner. For simplicity, we might assume that flux movement occurs with equal amplitude in all directions including between adjacent and diagonal pixels. We can describe this by summing the covariance terms corresponding to flux movement between each neighbouring pair of pixels together. Given these assumptions, we can describe the pixel kernel function of flux-conserved systematics for all pixels using:
@@
We note that we chose not to fit all the pixels on the detector, which means we ended up fitting pixels that have neighbours not included in our fit. As the flux may still ‘move between’ these pixels, we still account for the additional variance experienced by these border pixels by including these pixel fluxes in the sum of neighbouring fluxes, . However, there are no corresponding covariance terms included in the covariance matrix. The consequence of this is that the flux moving between pixels being fit and pixels not being fit will not be conserved; although this only affects pixels at the edges of our boundary, which will typically experience low flux. This effect can be reduced by including more pixels in the fit, although this comes at an increased computational cost.
@@ 2.4.2 Independent pixel systematics
Each pixel may also have systematics which are independent to neighbouring pixels, such as due to persistence from cosmic ray hits or unmasked bad pixels. For example, the pixel light curves from the second LHS 1140c eclipse appear to show a jump in flux in only a couple of pixels (as highlighted in Figure 1). For these systematics, we assume they can be modelled with the same kernel function in time as for flux-conserving systematics and have a pixel kernel function given by:
where we refer to these as independent pixel systematics (IPS) with height scale, . Since the amplitude of systematics may vary significantly between pixels (e.g. due to persistence effects from a cosmic ray hitting specific pixels), we may choose to fit for the amplitude, , separately for each pixel. Checking which pixels recover large values of could help identify pixels exhibiting increased systematics.
@@ 2.4.3 Common systematics between all pixels
We may also encounter systematics which have the same shape in all pixels and are proportional to the flux on each pixel. For example, if the target star flares, that could produce the same systematic shape in all pixels. These systematics may therefore represent stellar activity or alternatively could result from variations in the throughput of the detector which are correlated in time. We note that these systematics should be similar to fitting the aperture extracted light curve with a GP as they are not conserved and each pixel is treated similarly. Our pixel kernel function for these systematics is described by:
We refer to these as common systematics (CS) with height scale . The limitations of our GP optimisation require these systematics to also share the same time kernel function as the flux-conserved systematics and independent pixel systematics. A comparison between the three previous systematics described is shown in Figure 2.
@@ 2.4.4 Interpixel capacitance (IPC)
This effect can be modelled by taking the image we expect to receive (including Poisson noise) and using a 3x3 convolution kernel, (not to be confused with a GP kernel function), which blurs the signal received in each pixel into its neighbouring eight pixels. The kernel will always sum to unity which results in no change to the overall flux received. We can parameterise this convolution kernel using Equation 11 with the coupling between neighbouring pixels to the left or right, for pixels above and below and for diagonal pixels. More complex coupling is also possible but this parameterisation is common in the literature (Moore et al. 2006; Morrison et al. 2023) as follows:
Morrison et al. (2023) suggested that for the MIRI detector while is significantly smaller yet still non-zero. The exact extent of IPC is difficult to quantify and is claimed to vary between observations so we rely on fitting for and for each observation.
We need to account for both the amplitude of white noise for each pixel and the effect of IPC in our combined pixel kernel function . However, as IPC can be modelled using linear transformations, it is more natural to describe this by the covariance matrix generated by this kernel function, . Suppose each pixel observes Poisson noise with variance - which in the absence of IPC is independent between pixels. This can be described by a diagonal covariance matrix, , which contains the variance of each pixel along its diagonal (the pixels must be sorted into some arbitrary order). We can simulate IPC by taking a random draw from and then convolving with our IPC kernel from Equation 11. Note we can construct a matrix which performs this convolution for each pixel via the matrix multiplication: , where is our random draw of white noise with the effect of IPC simulated. The covariance matrix of is then given by:
@@
2.4.5 Background flux and systematics along rows and columns
@@
where is the amplitude of overall background scatter, is the amplitude of correlated scatter in each row and similarly for each column. By fitting for the amplitude of , and , we can constrain how much remaining background scatter is left by the background subtraction. We could potentially use this to compare different background subtraction procedures. We note that this does not give us the ability to detect whether the background has been over-subtracted or under-subtracted. Instead, it simply allows us to see whether there is a lot of correlated scatter between rows, columns, or all pixels.
@@ 3 Testing on simulated data
To better examine how the pixel-fitting method compares to aperture photometry, we generated synthetic observations for which the true eclipse depth was known. This then allows us to examine how precisely and robustly each method recovers the eclipse depth in different circumstances. We examined how well each method performed when there were only flux-conserved systematics present, when systematics were added in only a single pixel light curve, and when there were systematics present across all pixels with the same shape (e.g. from a flare).